# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v1

@@ IV-C Manipulation under Imperfect Priors
Figure 9: Manipulation under obsolescent priors. Tac-Man demonstrates exceptional performance in scenarios characterized by obsolescent priors, often resulting from unpredictable perturbations that modify object kinematics. Such perturbations are common in dynamic, human-centric environments, frequently emerging from unanticipated human interactions. To assess Tac-Man’s resilience against these perturbations, we introduced random forces to one side of the object in scenarios (a) and (b), effectively altering the cabinet’s rotation axis in (a) and the drawer’s opening direction in (b). Despite the introduction of errors through these human-induced disturbances, Tac-Man adeptly adjusts, preserving the correct interaction direction and ensuring successful manipulation outcomes. For video demonstrations, refer to Supplementary Video S4.
@@ V Large-scale Verification in Simulation
Acknowledging the limitation of current datasets of objects with simpler joint types, we further extend our verification to include articulated objects designed with complex manipulation trajectories. These objects are uniquely characterized by trajectories that follow randomly sampled curves, introducing additional challenges to the manipulation task. In Sec. V-B, we elaborate on the preparation of these objects and present both qualitative and quantitative findings from our simulations. The results compellingly illustrate Tac-Man’s adeptness at navigating objects with a broad range of articulations, underscoring its effective adaptability and generalization across diverse scenarios.
@@ V-A Environmental Setup
Our simulation environment, depicted in Fig. 10, features a robot arm equipped with a gripper poised to engage an object’s handle. Initially, the robot is programmed to follow a preliminary direction $R^{1}$, applying Tac-Man to articulate the object. A critical aspect of Tac-Man is its reliance on the compliance of silica gel material to accurately track contact between the gripper and the handle. To mimic this in a rigid body simulation environment, we employ specialized designs for contact measurement, detailed in Sec. -B.
The simulation divides test objects into two categories: standard objects with prismatic or revolute joints from existing datasets, and randomly created custom objects with complex trajectories. From the PartNet-Mobility dataset [36], we select objects with detailed articulation annotations provided by GAPartNet [38], as outlined in Sec. -C. For the evaluation, 25 objects with prismatic joints and another 25 with revolute joints were pre-processed. A manipulation is deemed successful if a revolute joint rotates beyond $60.0^{\circ}$or a prismatic joint extends over $250.0~{}\mathrm{mm}$. Fig. 10(a) showcases a subset of these objects and their initial grasps.
To assess Tac-Man’s adaptability beyond basic joint types, we also evaluate it on articulated objects designed with randomly generated trajectories. These are constructed by sampling Bézier curve segments of various orders (2, 3, 4, and 5) on a 2D plane, ensuring no self-intersections. Each curve inspires a 3D playboard that defines a manipulation path, with a toy train at the start point acting as the handle. Physical collisions are configured to restrict the handle’s movement along the trajectory, with the construction process detailed in Sec. -D. Fig. 10(b) illustrates an example of a 3-order Bézier curve, its corresponding play-board, and the simulation’s initial state. Additional examples of generated play-boards are presented in Fig. 10(c), highlighting the increased complexity and manipulation challenge with higher-order curves. For the evaluation, 25 curves of orders 3 to 6 are sampled, each tested once. Success is achieved when the toy train is moved from one end of the curve to the other by the gripper.
V-B Results
@@
Figure 11: Examples of simulated manipulation. Each row captures a sequence of manipulating an articulated object, showcasing scenarios with basic joint types (a-c) and intricate trajectories (d-e). For each scenario, the depiction includes the initial and final states, the execution-recovery iterations, and the tactile patterns observed. This visual representation underscores the effectiveness of our simulation setup in demonstrating Tac-Man’s capability to adapt and succeed across varying articulation challenges. For video demonstrations, refer to Supplementary Video S5.
@@
Quantitative evaluations for objects with prismatic and revolute joints are presented in Tab. II (columns 2-3), showcasing a $100~{}\%$ success rate across all 50 objects (25 per joint type). Additionally, results for objects following intricate manipulation trajectories are tabulated in Tab. II (columns 4-7), with PB-$n$ denoting performance on trajectories of order $n$. Remarkably, Tac-Man attains a $100~{}\%$ success rate in all 200 trials (25 curves × 4 orders × 2 endpoints), underlining its proficiency and reliability in handling a broad spectrum of articulation patterns on a large scale without reliance on specific priors.
TABLE II: Simulation quantitative results
Category Pri. and Rev. Intricate Trajectories
Pri. Rev. PB-2 PB-3 PB-4 PB-5
Succ. (%) 100 100 100 100 100 100
@@ VI Discussion
Figure 12: Manipulation under misaligned preliminary direction. Even given a misaligned preliminary direction, Tac-Man is still able to use it to successfully open the drawer.
@@ VI-A Advancing Robotic Manipulation with Tac-Man
Our study has validated the efficacy of Tac-Man, a tactile-informed, prior-free approach for manipulating articulated objects. By leveraging tactile feedback, Tac-Man discerns discrepancies between the preliminary direction and actual interaction directions, allowing dynamic adjustments to ensure precise manipulations. A standout feature of Tac-Man is its operational independence from pre-existing knowledge about object kinematics, presenting a considerable advantage over traditional methods reliant on priors. This capability is particularly beneficial in environments where priors may be ambiguous (Sec. IV-B), imperfect (Sec. IV-C), unknown (Sec. IV-D), or obsolescent (Sec. IV-E)—conditions prevalent in dynamic, human-centric settings.
Tac-Man enhances robotic autonomy by obviating the need for manually inputting accurate object kinematic models, a common requirement in earlier strategies [3, 4, 5]. It sidesteps the limitations of methods that depend on visual perception for determining object kinematics [9, 10, 11, 13, 14, 15, 20], which are vulnerable to the ambiguities and inaccuracies intrinsic to visual data [39]. Unlike approaches that utilize multi-frame observations to clarify ambiguities [7, 12, 19, 18, 16, 17, 6, 8], Tac-Man does not presuppose the existence of specific articulation types, thus demonstrating its versatility in managing novel articulation mechanisms as evidenced in Sec. IV-D.
Moreover, Tac-Man circumvents the exhaustive data collection process typically required by learning-based methods to tackle articulated object manipulation challenges [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. The diversity and complexity of articulated objects, ranging from basic prismatic and revolute joints to more intricate mechanisms featuring simultaneous translation and rotation, pose a significant challenge for data-driven approaches. Yet, Tac-Man adeptly addresses these challenges without relying on extensive datasets, thereby facilitating autonomous exploration and data collection for previously unseen objects. This autonomous data-gathering capability enhances the data collection process’s versatility and substantially reduces human labor.
@@ VI-B Limitations and Future Work
Preliminary direction acquisition
The efficacy of Tac-Man hinges on the availability of a preliminary direction for manipulating an articulated object. This foundational assumption is supported by advances in research, such as the work by Mo et al. [71], which aims to identify critical interaction points and predict preliminary directions for interaction, like pulling or pushing, for engaging with articulated objects. Although these directions might not always be accurate, Tac-Man is specifically engineered to accommodate and correct such preliminary direction, effectively handling any associated uncertainties. To showcase Tac-Man’s resilience in this context, we experiment with a drawer, providing the robot arm with two intentionally erroneous directions. Despite these deviations, as depicted in Fig. 12, Tac-Man demonstrates its robust capability to adapt and successfully complete the drawer opening task.
This exploration into the reliance on preliminary direction underscores a limitation that also opens avenues for future research. Enhancing the accuracy of these directions or developing mechanisms within Tac-Man to autonomously generate and refine interaction directions could further improve its applicability and success rate across a wider array of articulated objects and scenarios. Investigating these enhancements would be a valuable direction for future work, potentially leading to even more adaptable and autonomous robotic manipulation systems.
@@
While Tac-Man demonstrates notable capabilities in the prior-free manipulation of articulated objects, an area for potential enhancement is its time efficiency. As indicated in Tab. III, the completion time for manipulation tasks using Tac-Man tends to exceed that of human performance. This disparity is largely attributed to the current implementation of Tac-Man, which operates without the benefits of closed-loop control.
TABLE III: Time for completing various manipulation tasks.
Case Fig. 6(e) Fig. 6(f) Fig. 6(g) Fig. 6(h) Fig. 7(b) Fig. 8(a) Fig. 8(b)
Time ($\mathrm{s}$) $\approx 480$ $\approx 600$ $\approx 480$ $\approx 180$ $\approx 750$ $\approx 280$ $\approx 900$
In its present form, Tac-Man employs a cautious, step-by-step exploratory approach. This method, devoid of pre-established object priors or learning, requires careful evaluation of, and response to, tactile feedback. The task duration figures reported in Tab. III primarily reflect this initial phase of exploration, critical for the robot to accurately interpret and interact with the object.
@@
Looking ahead, enhancing Tac-Man to include closed-loop control elements could significantly improve its time efficiency. Such a development would enable quicker and more responsive adaptations during initial exploration, potentially reducing the overall time required for manipulations. It would lead to a more balanced approach, achieving manipulation robustness and efficiency on a par with human performance, especially in dynamic environments as discussed in Sec. IV-E.

### v3

@@ IV-D Manipulation under Unknown Priors
Figure 9: Manipulation under obsolescent priors. Tac-Man demonstrates exceptional performance in scenarios characterized by obsolescent priors, often resulting from unpredictable perturbations that modify object kinematics. Such perturbations are common in dynamic, human-centric environments, frequently emerging from unanticipated human interactions. To assess Tac-Man’s resilience against these perturbations, we introduced random forces to one side of the object in scenarios (a) and (b), effectively altering the cabinet’s rotation axis in (a) and the drawer’s opening direction in (b). Despite the introduction of errors through these human-induced disturbances, Tac-Man adeptly adjusts, preserving the correct interaction direction and ensuring successful manipulation outcomes. For video demonstrations, refer to Supp Video S4.
@@ V Large-scale Verification in Simulation
Acknowledging the limitation of current datasets of objects with simpler joint types, we further extend our verification to include articulated objects designed with complex manipulation trajectories. These objects are uniquely characterized by trajectories that follow randomly sampled curves, introducing additional challenges to the manipulation task.
To showcase the effectiveness of Tac-Man, we select several representative prior-based and prior-free methods as baselines, as detailed in Sec. V-B. In Sec. V-C, we present both qualitative and quantitative results from our simulations. The findings compellingly demonstrate Tac-Man’s proficiency in manipulating objects with a wide range of articulations, highlighting its adaptability and superior generalization across diverse scenarios compared to the baseline methods.
@@ V-A Environmental Setup
Our simulation environment, depicted in Fig. 10, features a robot arm equipped with a gripper poised to engage an object’s handle. Initially, the robot is programmed to follow a preliminary direction $\boldsymbol{R}^{1}$, applying Tac-Man to articulate the object. A critical aspect of Tac-Man is its reliance on the compliance of silica gel material to accurately track contact between the gripper and the handle. To mimic this in a rigid body simulation environment, we employ specialized designs for contact measurement, detailed in Sec. -A.
The simulation divides test objects into two categories: standard objects with prismatic or revolute joints featuring various handle designs from existing datasets, and randomly created custom objects with complex trajectories. From the PartNet-Mobility dataset [36], we select objects with detailed articulation annotations provided by GAPartNet [38], as outlined in Sec. -B. For the evaluation, $25$ objects with prismatic joints and another $25$ with revolute joints were pre-processed. A manipulation is deemed successful if a revolute joint rotates beyond $60.0^{\circ}$or a prismatic joint extends over $250.0~{}\mathrm{mm}$. Fig. 10(a) showcases a subset of these objects and their initial grasps.
To assess Tac-Man’s adaptability beyond basic joint types, we also evaluate it on articulated objects designed with randomly generated trajectories. These are constructed by sampling Bézier curve segments of various orders ($2$, $3$, $4$, and $5$) on a $2$D plane, ensuring no self-intersections. Each curve inspires a $3$D playboard that defines a manipulation path, with a toy train at the start point acting as the handle. Physical collisions are configured to restrict the handle’s movement along the trajectory, with the construction process detailed in Sec. -C. Fig. 10(b) illustrates an example of a $3$-order Bézier curve, its corresponding play-board, and the simulation’s initial state. Additional examples of generated play-boards are presented in Fig. 10(c), highlighting the increased complexity and manipulation challenge with higher-order curves. For the evaluation, $25$ curves of orders $2$ to $5$ are sampled, each tested once. Success is achieved when the gripper moves the toy train from one end of the curve to the other.
V-B Selected Baselines
To demonstrate the effectiveness of Tac-Man, we also select several representative prior-based or prior-free methods designed for the manipulation of articulated objects to conduct comparisons with Tac-Man. Specifically, we evaluate the performance of three distinct techniques: the prior-based pre-planned trajectory generation method, impedance control in a prior-free setting, and a vision-based learning approach, FlowBot3D [19]. Below, we detail the setup for each of these methods:
Prior-based pre-planned trajectory
This method is highly effective when the kinematic priors of articulated objects are precisely known. However, meeting this requirement proves challenging in the real world due to pervasive uncertainties and unexpected perturbations [39]. To account for this in our simulations, we introduce noise $\xi$ into the ideal kinematic model of the test objects. This noise $\xi$ is modeled as a normal distribution with a mean of zero and a variance of $0.01$. We explore the impact of this noise on the pre-planned trajectory generation method under the same initial conditions as Tac-Man for different joint types. For revolute joints, the relationship between the actual handle trajectory radius $r_{gt}$ and the manipulated trajectory radius $r_{m}$ (both in meters) is:
$r_{m}=r_{gt}+\xi.$ (25)
For prismatic joints, the relationship between the actual handle trajectory angle $\theta_{gt}$ and the manipulated trajectory angle $\theta_{m}$ (both in radians) is:
$\theta_{m}=\theta_{gt}+\xi.$ (26)
For Bézier curves, the relationship between the actual control point position $\boldsymbol{p}_{gt}$ and the manipulated trajectory control point position $\boldsymbol{p}_{m}$ (both in meters), excluding the starting point, is:
$\boldsymbol{p}_{m}=\boldsymbol{p}_{gt}+\begin{bmatrix}\xi&\xi\end{bmatrix}^{\mathsf{T}}.$ (27)
Impedance control
Impedance control, introduced by Hogan [81], is a robust technique widely used in robotic manipulation tasks. We compare impedance control with Tac-Man, following the same preliminary direction. In this context, impedance control serves as a prior-free approach for manipulating articulated objects and serves as a baseline in an ablation study to assess the essential role of tactile sensing in the Tac-Man.
FlowBot3D
FlowBot3D, introduced by Eisner et al. [19], is a recent advancement using modern learning techniques for generalized manipulation based on visual feedback. It employs a flow prediction model to predict part mobility and compute actions accordingly. It should be noticed that, to align with our simulation settings, we apply only the flow prediction model and action computation strategy. We assess its performance using the same tasks described in Sec. V-A as those used for Tac-Man, without additional fine-tuning.
V-C Results
@@
Figure 11: Examples of simulated manipulation. Each row captures a sequence of manipulating an articulated object, showcasing scenarios with basic joint types (a-c) and intricate trajectories (d-e). For each scenario, the depiction includes the initial and final states, the execution-recovery iterations, and the tactile patterns observed. This underscores the effectiveness of our simulation setup in demonstrating Tac-Man’s capability to adapt and succeed across varying articulation challenges. For video demonstrations, refer to Supp Video S5.
@@
Figure 12: Barplot of the distance-weighted success rates of different methods across six articulated object types. Tac-Man achieves $100\%$ success rate, outperforming the other three baseline methods across all the articulated object types ( Tested on Isaac Sim using our settings without fine-tuning on PB-$n$).
The quantitative results for the three baseline methods and Tac-Man are summarized in Fig. 12. To provide a more informed measure of performance beyond binary success or failure, we introduce a distance-weighted success rate $SR_{W}$, defined as:
$SR_{W}=\frac{\min(d_{A},d_{E})}{d_{E}}\times 100\%,$ (28)
where $d_{E}$ represents the expected contact point moving distance required for success, as specified in Sec. V-A, and $d_{A}$ is the actual handle moving distance calculated from the final handle position in each trial.
Fig. 12 presents both mean and standard deviation of the weighted success rate across objects under different categories. Specifically, columns $1$-$2$ display weighted success rates for manipulating objects with prismatic and revolute joints, while columns $3$-$6$ show performance on intricate manipulation trajectories, with PB-$n$ indicating performance on Bézier trajectories of order $n$. Notably, Tac-Man achieves a $100\%$ weighted success rate across all objects with various handle designs, markedly outperforming the three baseline methods and verifying its effectiveness. Key observations for the baseline methods include:
  • Pre-planned trajectory methods: In scenarios lacking tactile feedback, these methods are vulnerable to noise, affecting their performance.
  • Impedance control: Following the same preliminary direction as Tac-Man, yet without object kinematics priors and tactile feedback, this method proves effective for objects with prismatic joints. However, objects requiring movements beyond a straight trajectory may quickly surpass the compliance limits, resulting in failures.
  • FlowBot3D: This method demonstrates effectiveness with familiar objects featuring prismatic and revolute joints. However, it underperforms on unfamiliar, intricate trajectories.
@@ VI Discussion
Figure 13: Manipulation under imperfect initial grasp and misaligned preliminary direction. Even given an imperfect initial grasp and a misaligned preliminary direction, Tac-Man is still able to use it to successfully open the drawer.
@@ VI-A Advancing Robotic Manipulation with Tac-Man
Our study has validated the efficacy of Tac-Man, a tactile-informed, prior-free approach for manipulating articulated objects. The core of Tac-Man is leveraging tactile feedback to discern discrepancies between the preliminary direction and actual interaction directions, informing dynamic adjustments to ensure precise manipulations. A standout feature of Tac-Man is its operational independence from pre-existing knowledge about object kinematics, presenting a considerable advantage over traditional methods reliant on priors. This capability is particularly beneficial in environments where priors may be ambiguous (Sec. IV-B), imperfect (Sec. IV-C), unknown (Sec. IV-D), or obsolescent (Sec. IV-E)—conditions prevalent in dynamic, human-centric settings. While classical control methods like impedance control can operate without object priors, they may fail with complex trajectories, as shown in Fig. 12.
Tac-Man enhances robotic autonomy by obviating the need for manually inputting accurate object kinematic models, a common requirement in earlier strategies [4, 5, 6]. It sidesteps the limitations of methods that depend on visual perception for determining object kinematics [10, 11, 12, 14, 15, 16, 21], which are vulnerable to the ambiguities and inaccuracies intrinsic to visual data [39], as demonstrated by the reduced effectiveness of pre-planned methods under noise in Fig. 12. Unlike approaches that utilize multi-frame observations to clarify ambiguities [8, 13, 20, 19, 17, 18, 7, 9], Tac-Man does not presuppose the existence of specific articulation types, thus demonstrating its versatility in managing novel articulation mechanisms as evidenced in Sec. IV-D and Fig. 12.
Moreover, Tac-Man circumvents the exhaustive data collection process typically required by learning-based methods to tackle articulated object manipulation challenges [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. The diversity and complexity of articulated objects, ranging from basic prismatic and revolute joints to more intricate mechanisms featuring simultaneous translation and rotation, pose a significant challenge for data-driven approaches, as shown by FlowBot3D’s lower success rates on unseen, intricate trajectories. Yet, Tac-Man adeptly addresses these challenges without relying on extensive datasets, thereby facilitating autonomous exploration and data collection for previously unseen objects. This autonomous data-gathering capability enhances the data collection process’s versatility and substantially reduces human labor.
@@ VI-B Limitations and Future Work
Compatibility across different sensors
Tac-Man’s effectiveness primarily stems from its ability to compensate for contact deviation, which is modeled in a geometric perspective so that both direction and magnitude of deviation can be obtained, as detailed in Sec. III-E. This paper demonstrates the implementation of Tac-Man with GelSight-based sensor, as one economic representative for those sensors with the ability to measure contact geometry. We would like to clarify that the high-resolution property of GelSight is not a major consideration in this work, as indicated in the setup for experiments. Actually, all these sensors can be effectively adopted within the Tac-Man framework and should have a very similar formulation, with only minor variation in the pre-processing procedure where contact deviation geometry is extracted from raw sensor data. For example, the TacTip family [63] uses an array of pins to detect surface deformations, which can be mapped to contact points in a manner similar to our current implementation. Therefore, one important future work will be to investigate whether the extraction of contact deviation geometry with different sensors can be generalized to some degree to improve the succinctness of Tac-Man formulation.
Force/torque sensors, while primarily measuring forces and torques, can also be adapted for use within the Tac-Man framework, with an extra step of calculating geometric contact deviation, and thus the magnitude for compensation, from the force/torque measurements by Hooke’s Law [66]:
[…]
where $\boldsymbol{c}\in\mathbb{R}^{6}$ represents the geometric contact deviation, and $\boldsymbol{f}\in\mathbb{R}^{6}$ the force/torque measurement. $K$ refers to the contact stiffness matrix, which can be constant or varying depending on the interaction surface, and often needs proper calibration before use. Therefore, another interesting future direction will be the investigation of contact stiffness calibration, which may further motivate the possibility of exploring force-based formulation for Tac-Man, where obtaining the magnitude of contact deviation from force becomes a major point.
Therefore, Tac-Man in principle has the ability to accommodate various types of sensors. Conducting comparative studies to evaluate the impact of different sensor types on Tac-Man’s performance in different scenarios remains an important area for future work. These investigations could assist in optimizing sensor selection under specific tasks and potentially lead to a more generalized and succinct formulation of Tac-Man that can seamlessly incorporate data from diverse sensor types.
Initial grasp generation and preliminary direction acquisition
The effectiveness of Tac-Man relies on obtaining a stable initial contact $\mathcal{C}_{1}$ and having a preliminary (rough) direction to manipulate an articulated object. In this study, we manually establish the initial (rough) grasp and preliminary direction for all experiments. Crucially, Tac-Man’s success stems from its ability to adjust control input on the fly based on tactile feedback; hence, this initial grasp need not be precise and can sometimes even be intentionally wrong. To demonstrate Tac-Man’s resilience, we conducted experiments on a drawer (see Fig. 13), providing the robot arm with intentionally incorrect grasps and directions. Despite these inaccuracies, Tac-Man adapts and successfully completes the task of opening the drawer.
Although Tac-Man does not require prior knowledge, such as object kinematics and learning policy, which was once deemed essential in the manipulation of articulated objects, its dependence on initial grasp and preliminary direction highlights a limitation but also suggests a potential area for future research. Improving the precision of learning-based initial grasps and directions [82, 45, 83], or developing capabilities within Tac-Man to autonomously refine them, could enhance its effectiveness and broaden its application across various scenarios involving articulated objects. These improvements could potentially lead to more adaptable and autonomous robotic manipulation systems.
Figure 14: Impact of grasping force. Extreme forces significantly reduce the efficiency of Tac-Man by affecting the computation of $y$, the maximum distance during the execution stage, in the optimization program (7). This impact results in an increase in both the number of (a) execution-recovery cycles and (b) execution times. Crucially, within the elastic deformation region, Tac-Man yields successful manipulations of articulated objects despite varied execution time, demonstrating Tac-Man’s robustness.
Grasping force
The initial grasping force critically influences Tac-Man’s efficiency by directly impacting the stable contact state $\mathcal{C}_{1}$ and associated tactile feedback. While it minimally affects the computation of the optimal transformation matrix in Eq. 23 (which relies on relative changes to $\mathcal{C}_{1}$ for contact stabilization), it significantly influences the calculation of $t$, the maximum distance during the execution stage. This $t$ value, derived from the optimization program (7), depends on absolute changes. Simulation results in Fig. 14 demonstrate that both excessively large and small grasping forces prolong manipulation times. This aligns with intuition: larger forces may exceed the gripper’s maximum elastic deformation capacity, while smaller forces increase slippage risk. Both scenarios are more likely to violate the stable contact constraint (2), reducing the maximum $t$ value from program (7) and necessitating additional iterations between execution and recovery stages. Crucially, within the elastic deformation region, Tac-Man yields successful manipulations of articulated objects despite varied execution time, demonstrating its robustness. Based on these findings, we practically select a grasping force that yields a deformation of approximately half the maximum normal elastic deformation $e_{n}$. While this approach is effective, future work should explore methods for adaptively selecting the optimal grasping force to further enhance performance across various manipulation scenarios.
@@
While the Tac-Man demonstrates notable capabilities in the prior-free manipulation of articulated objects, an area for potential enhancement is its time efficiency. As indicated in Tab. II and Fig. 15, the completion time for manipulation tasks using Tac-Man tends to exceed that of human performance. This disparity is largely attributed to the current implementation of Tac-Man, which operates without the benefits of closed-loop control.
TABLE II: Time in second for completing tasks in the real world.
Task Fig. 6(e) Fig. 6(f) Fig. 6(g) Fig. 6(h) Fig. 7(b) Fig. 8(a) Fig. 8(b)
Time (s) $\approx 480$ $\approx 600$ $\approx 480$ $\approx 180$ $\approx 750$ $\approx 280$ $\approx 900$
Figure 15: Time in second for completing tasks in simulation studies.
In its present form, Tac-Man employs a cautious, step-by-step exploratory approach. This method, devoid of pre-established object priors or learning, requires careful evaluation of, and response to, tactile feedback. The task duration figures reported in Tab. II and Fig. 15 primarily reflect this initial phase of exploration, critical for the robot to accurately interpret and interact with the object.
@@
Looking ahead, enhancing Tac-Man to include velocity-based control elements could significantly improve its time efficiency. Such a development would enable quicker and more responsive adaptations during initial exploration, potentially reducing the overall time required for manipulations. It would lead to a more balanced approach, achieving manipulation robustness and efficiency on a par with human performance, especially in dynamic environments as discussed in Sec. IV-E.
Integrating Tac-Man into learnable framework
Tac-Man, as discussed in Sec. II-E, is a model-based tactile servoing method. This approach offers benefits such as reduced data collection requirements and enhanced explainability [84]. However, integrating Tac-Man into a learnable framework represents a promising avenue for future research. This integration aims to generalize the strategy of utilizing tactile sensing in articulated object manipulation to a broader range of challenges, including dexterous manipulation [3], cable following [74], surface tracking [53], and others.